please help me solve this problem by substitution
(x-4y+z=6 & 2x+5y-z=7 & 2x-y-z=1)
What would x, y & z equal?

Answer by stanbon(75887)

x-4y+z =6
2x+5y-z=7
2x-y-z =1
------------------
Use any method you know to get:
x = 4
y = 1
z = 6
=============
Cheers,
Stan H.
=============